OpenStudy (anonymous):

Solve the system of linear equations below. x + y = 4 2x + 3y = 0

OpenStudy (anonymous):

Hi Welcome to Openstudy!! Are you familiar with the elimination or substitution method?

OpenStudy (anonymous):

it has to be negative

OpenStudy (goformit100):

Welcome to Open Study .................. Sir/Ma'am @caleb_gajewsky

OpenStudy (anonymous):

it's ma'am, and thank you

OpenStudy (muzzack):

excuse me @austinisawesome12 what do you mean it has to be negative?

OpenStudy (anonymous):

I am not aware

OpenStudy (muzzack):

x + y = 4 we have to isolate y in this equation y=-x+4

OpenStudy (anonymous):

oh yeah true

OpenStudy (muzzack):

then we replace y in this equation: 2x + 3y = 0 2x+3(-x+4)=0

OpenStudy (muzzack):

now we solve for x

OpenStudy (anonymous):

-2x = 12 x = -6 ?

OpenStudy (muzzack):

2x+3(-x+4)=0 ------- 2x-3x+12=0 ------- -x+12=0 ------- -x=-12 ------- x=12

OpenStudy (muzzack):

now we replace x with the 2nd equation to solve for y 2(12) + 3y = 0 ----- 24+3y=0 3y=-24 y=-8

OpenStudy (muzzack):

now you have both answers x=12 and y=-8

OpenStudy (anonymous):

Using the elimination method, you will eliminate one of the variables in order to solve for the other variable. To do this: 1. Decide which variable you want to eliminate, say for example you want to solve for x. 2. To eliminate x, you have to multiply the first equation by -2. 3. Then, add both the equations. Simplify it and solve for y. 4. Once you already solved for y, substitute this y-value into on of the equations. 5. From that, you can be able to solve for x. Substitution method is like what Muzzack is doing. Hope this helps! Enjoy! (^_^)

OpenStudy (anonymous):

thank y'all !